OpenStudy (anonymous):

A farmer wants to fence off a 22,500 square feet area in a rectangular field and then divide it in half with a fence parallel to one of the sides of the rectangle. How can the farmer do this so as to minimize the cost of the fence?

myininaya (myininaya):

if we draw a line to divide the rectangle in half wouldn't it be parallel to two sides and not just one of the sides of the rectangle?

OpenStudy (anonymous):

X*y=22500 s = 2x + 2y + x Y= 22500/x Min(3x +45000/x)

OpenStudy (anonymous):

min{3 x+45000/x} = 300 sqrt(6) at x = 50 sqrt(6)

myininaya (myininaya):

i don't know about this question i'm still confused about that one part

OpenStudy (anonymous):

so what would the final answer be?

OpenStudy (radar):

Continue on with ivanbaj data: x*y=22,500 sq ft s (fence length) will be 2x + 2y +x (the extra x is the divider portion) Also y in terms of x is 22,500/x (from the area formula) substitute for the y in the amount of fence formula. getting s=2x + 2(22500/x) + x s=3x + 45000/x now differentiate ds/dx= 3-45000x^-2 Set to 0 to find minima \[3-45000/x ^{2}=0\] multiply thru by x^2 getting \[3x ^{2}-45000=0\]\[3x ^{2}=45000\] \[x ^{2}=15000\]\[x=\sqrt{15000}=\sqrt{2500*6}=50\sqrt{6}\] \[y=22500\div 50\sqrt{6}=75\sqrt{6}\] The cost of the fencing will be 300 times sq rt of 6 times what ever the fence cost per foot. Get out of the water and practice these problems with the intent to learn procedures not 'just get answers"

myininaya (myininaya):

radar i'm still confused about the question? how can draw a line parallel to one of the sides of the rectangle without drawing a line parallel to two sides of the rectangle

myininaya (myininaya):

it doesn;t make sense to me lol

OpenStudy (radar):

It would be parallel to two of the sides and perpendicular to two of the sides.

myininaya (myininaya):

right but why does it say draw it parallel to one of the sides?

OpenStudy (radar):

That is to assure that you won't divide the rectangle by a diagonal. Noticed it didn't say that it would not be parallel to the other side.

myininaya (myininaya):

i guess you are right lol

OpenStudy (radar):

Yes, sometimes the instructions are more complex than the problem lol

myininaya (myininaya):

i think it would be fun to the diagonal problem lol

OpenStudy (radar):

Also English can sometimes lead to part of the problem (by the way it is used)

myininaya (myininaya):

yes

OpenStudy (radar):

Anyway I hope that the curiousmathstudent understands it. Have a nice day myininaya, and I have noticed you providing lots of good help. bye

myininaya (myininaya):

you too you're awesome too!